OpenStudy (anonymous):

if an arrays first 6 steps are: 1,2,3,4,5,6 then what will be the sum of all the values between 100th and 1th steps of array?

OpenStudy (anonymous):

if it was 6,, it would be 1 + 2 + 3 + 4 + 5 + 6= 21 if it was 5 ,, it would be 1 + 2 + 3 + 4 + 5 = 15 if it was 4,, it would be 1 + 2 + 3 + 4= 10 if it was 3,, it would be 1 + 2 + 3 = 6 then u can see that in first it rises 4 and then 5 and then 6 so why don't we use this?

OpenStudy (anonymous):

for 2 its 1 + 2 = 3 for 1 its 1 so it always + 1,, + 2,, + 3,, + 4 so its n+N i think

OpenStudy (angela210793):

hmm i thought 1.......100 and...1+100=101 also 2+99=101 so u can do 101*50 (or 49...) not really sure how many couples like this r formed

OpenStudy (nowhereman):

so what is N?

OpenStudy (anonymous):

its the value of the raise :d between the sums :D the value of the raise of sum always raises 1

OpenStudy (anonymous):

but n+N won't work..

OpenStudy (nowhereman):

Try to do it angelas (actually Gauß's) way!

OpenStudy (anonymous):

:D oke

OpenStudy (anonymous):

but can't we solve this on my way too?

OpenStudy (anonymous):

to use the value of the raise

OpenStudy (angela210793):

i didn't really get ur way...

OpenStudy (nowhereman):

What you do is nothing else than reformulating the inductive formula, which will not lead to an explicat one.

OpenStudy (anonymous):

ok ^^

OpenStudy (anonymous):

nowhereman isn2T angela better then me in maths ? she says she is not but i say she is better then me in maths.

OpenStudy (nowhereman):

That's too general to say.